OpenStudy (anonymous):

Which of the following is a polynomial with roots 2, –3i, and 3i ?

OpenStudy (anonymous):

options: x3 + 4x2 + 9x + 24 x3 – 4x2 + 9x – 24 x3 + 2x2 + 9x + 18 x3 – 2x2 + 9x – 18

OpenStudy (anonymous):

help please?

OpenStudy (anonymous):

(x-2)(x+3i)(x-3i)

OpenStudy (anonymous):

expand the equation

OpenStudy (anonymous):

(x-2)(x^2+9)

OpenStudy (anonymous):

can you do the rest?

OpenStudy (anonymous):

I've been having a lot of trouble with it, would you mind working it out for me?

OpenStudy (anonymous):

Well sure

OpenStudy (anonymous):

The first step is given by the equation

OpenStudy (anonymous):

(x-2)(x+3i)(x-3i)

OpenStudy (anonymous):

Those are the roots of the equation... The answers in the parenthesis will get you 0 if you put in the 2,3i and -3i for x.

OpenStudy (anonymous):

then when multiplying it out to get the equation i did the part with 3i and -3i first

OpenStudy (anonymous):

(x+3i)(x-3i) = x^2 - 3xi + 3xi -9i^2

OpenStudy (anonymous):

i^2 = -1

OpenStudy (anonymous):

and the terms in the middle cancel out

OpenStudy (anonymous):

so that gets you x^2 + 9

OpenStudy (anonymous):

then you multiply that with (x-2)

OpenStudy (anonymous):

(x-2)(x^2+9)= x^3 + 9x - 2x^2 -18

OpenStudy (anonymous):

rearrange the terms and there is your answer

OpenStudy (anonymous):

Any questions?

OpenStudy (anonymous):

ohhh so it's the last one right? now it makes sense to me

OpenStudy (anonymous):

No problem. =D